OpenStudy (anonymous):

integrate by partial fractions, not sure,. i want to integrate 1/((x^2*(x^2-15)) should i do partial fraction 1/((x^2*(x^2-15)) = A / x + B/x^2 + C/(x-sqrt 15) + D/(x+sqrt 15) , or 1/((x^2*(x^2-15)) = A / x + B/x^2 + Cx+d / ( x^2 - 15)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

either way worked out

OpenStudy (anonymous):

i checked my program

OpenStudy (anonymous):

\[1/15\int\limits_{}^{}(x ^{2}-(x ^{2}-15))dx/(x ^{2}(x ^{2}-15))\]\[1/15[\int\limits_{}^{}dx/(x ^{2}-15)-\int\limits_{}^{}dx/x ^{2}\]

OpenStudy (anonymous):

now go ahead.

OpenStudy (anonymous):

yes that works, but i want to use partial fractions

OpenStudy (anonymous):

1/(x^2(x^2-15)) = A/x + B/x^2 + (cx+d)/(x^2-15)

OpenStudy (anonymous):

thats some trick you got there :)

OpenStudy (anonymous):

partial fraction will be avoided when you have better choices...

OpenStudy (anonymous):

the tricks will help u and its very time saving...

OpenStudy (anonymous):

i guess was wondering, is it ok to have , a quadratic that is not reducible , ok let me state this correctly

OpenStudy (anonymous):

how did you know to seperate them like that

OpenStudy (anonymous):

only practice..........

OpenStudy (anonymous):

i can make it harder, one sec

OpenStudy (anonymous):

int (1/ (x^3 (x^2-15))

OpenStudy (anonymous):

ok thats a good partial fractions problem

OpenStudy (anonymous):

yea...

OpenStudy (anonymous):

here you can use partial fractions because no other way............

OpenStudy (anonymous):

ok, now we have two options it would seem, my question really boils down to this quadratic

OpenStudy (anonymous):

the quadratic is not irreducible,

OpenStudy (anonymous):

integrate by partial fractions, not sure,. i want to integrate 1/((x^2*(x^2-15)) should i do partial fraction 1/((x^2*(x^2-15)) = A / x + B/x^2 + C /x^3 +D/(x-sqrt 15) + e/(x+sqrt 15) , or 1/((x^2*(x^2-15)) = A / x + B/x^2 + Cx+d / ( x^2 - 15)

OpenStudy (anonymous):

here u can use the one which i gave first to reduce it . it would give you \[1/15[\int\limits_{}^{}dx/(x(x ^{2}-15)-\int\limits_{}^{}dx/x ^{3}]\]

OpenStudy (anonymous):

my question was, both approaches are legal?

OpenStudy (anonymous):

does that work, im not sure

OpenStudy (anonymous):

both are correct.

OpenStudy (anonymous):

oh ok, i thought the quadratic MUST be irreducible

OpenStudy (anonymous):

but when u try to find cx+d/(x^2-15) then may be again u have to do partial fractions...

OpenStudy (anonymous):

are you sure this is correct 1/15[∫dx/(x(x2−15)−∫dx/x3], i did it and got a different answer

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

[ x^3 - x(x^2-15) ] / (x^4 (x^2 - 15)

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

im trying to reverse your result, checking if its equivlanet

OpenStudy (anonymous):

((x^2)-(x^2-15))/(x^3(x^2-15))

OpenStudy (anonymous):

oh it works

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my math was off,

OpenStudy (anonymous):

again there is another way to do it just substitute x^2=y.

OpenStudy (anonymous):

thats not going to work, since 2x dx = dy

OpenStudy (anonymous):

it is again more easier ....

OpenStudy (anonymous):

then dx = dy / (2x )

OpenStudy (anonymous):

oh you mean for the first problem ?

OpenStudy (anonymous):

your integration is xdx/(x^4(x^2-15)

OpenStudy (anonymous):

for the second one...

OpenStudy (anonymous):

try it... it surely works..

OpenStudy (anonymous):

oh , but its not simple

OpenStudy (anonymous):

lets try it for the first integral

OpenStudy (anonymous):

oh ok , so s

OpenStudy (anonymous):

its not for the first one.... its only applicable for the second one....

OpenStudy (anonymous):

these some shortcut to approach....

OpenStudy (anonymous):

i have 2 go now.....

OpenStudy (anonymous):

dx = dy / ( 2 sqrt y )

OpenStudy (anonymous):

do you agree to that?

OpenStudy (anonymous):

have u understood properly?

OpenStudy (anonymous):

yes, i just want to make sure i understand your subsitution

OpenStudy (anonymous):

dx/(x^3(x^2-15) gives you dy/(2y^2(y-15))

OpenStudy (anonymous):

i get dy / ( 2 y^(3/2) (y-15)

OpenStudy (anonymous):

y=x^2 dy=2xdx dx/(x^3(x^2-15))= 2xdx/(2x^4(x^2-15)) = dy/(2y^2(y-15))

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

you have to solve for dx

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

check it , thats correct...

OpenStudy (anonymous):

2x dx = dy , so dx = dy/(2x)

OpenStudy (anonymous):

ok...let me to make u clear...

OpenStudy (anonymous):

oh, you multiplied top and bottom by 2x, one sec

OpenStudy (anonymous):

yesssssssssss

OpenStudy (anonymous):

be fast i have to go .... but its nice to solve maths with you?

OpenStudy (anonymous):

very cleverl

OpenStudy (anonymous):

thanks , bye

OpenStudy (anonymous):

sure we will talk again

OpenStudy (anonymous):

thats the magic of maths...........

OpenStudy (anonymous):

my way probably works, i have to work on it :)

OpenStudy (anonymous):

my substitution

OpenStudy (anonymous):

what do u do?

OpenStudy (anonymous):

i solve for dx =

OpenStudy (anonymous):

from where r u?

OpenStudy (anonymous):

i live in usa, near new york

OpenStudy (anonymous):

in which class are u?

OpenStudy (anonymous):

2x dx = dy , so dx = dy/(2x) , so dx = dy/ ( 2 sqrt y)

OpenStudy (anonymous):

but it comes out the same

OpenStudy (anonymous):

since x^3 = sqrt y ^3

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

are u a student in college or school?

OpenStudy (anonymous):

college

OpenStudy (anonymous):

yes thats right..

OpenStudy (anonymous):

very nice, both ways work

OpenStudy (anonymous):

ok bye.......

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

i like your way though

OpenStudy (anonymous):

it never occured to me you can multiply top and bottom by x